find the total thickness of two pices of wood that a carpenter glued together if one is 5/16 inch thick and the other 7/8 inch thick
Found 2 solutions by stanbon, dragonwalker:
Answer by stanbon(75887)

find the total thickness of two pices of wood that a carpenter glued together if one is 5/16 inch thick and the other 7/8 inch thick
------
(5/16) + (7/8)
---
= (5/16) + (14/16)
----
= 19/16
---
= 1 3/16
==========================================
Cheers,
Stan H.
===============

Answer by dragonwalker(73)

So simply put you want to add the fractions as follows:
5/16 + 7/8
Firstly we want to make the denominators the same (the bottom number of the fractions). This is also called a common denominator as both denominators are 'common' to each other.
So we have to decide what the two bottom numbers have in common. What is the smallest number that they will both fit into. The easiest way is to see if the smaller number(s) will fit into the largest number.
In this case the 8 would fit perfectly into 16 so we can use 16 without a problem.
The thing is we can't just put a 16 instead of the 8 because the fraction has to be the same. So in order to turn the 8 into 16 what are we doing to the 8? i.e. how are we changing it?
well 8 is half of 16 so in order to make it 16 we have to multiply it by 2.
So to keep the fraction the same we have to multiply the top by 2 as well so that it still is the same size in relation to the bottom number
So: 7/8 now becomes 14/16
Now we can add the fractions as so:
5/16 + 14/16
when adding fractions think of a pie. The pie is cut into sections, in this case one pie or whole number has 16 pieces.
The first fraction has 5 pieces and the second one has 14 pieces. If we wanted to know how many pies we had we would add how many pieces we had.
So we add just the top number as the bottom number just tells us how many of the pieces fit into one pie.
So 5/16 + 14/16 = 19/16
But look the top number has more pieces than the pie tin can fit. So we take 16 pieces away and fill one pie pan. How many do we have left:
19 - 16 = 3
So in this case we have one whole and then 3 pieces. The 3 pieces of 16 possible in a pie pan.
So we write 19/16 as 1 whole and then 3/16
= 1 3/16 !!!
So the thickness of the wood is 1 3/16 inches